# Speeding up an unsteady flow simulationM. HANEK, J. PAPEŽ, and J. ŠÍSTEK

### v1

@@ 6 Results
In this section, we present numerical results for the simulation of incompressible flow around a unit sphere. We focus on multiple variants of the three-level BDDC method for the Poisson problem for pressure corrector (3) with a fixed matrix and time-dependent right-hand side vector. The computational mesh of Taylor–Hood hexahedral elements leads to 1.4M unknowns for pressure, and it is decomposed into 1024 subdomains; see Fig. 1. The time step is set to 0.05 s.
@@
The problems for velocity components (2) are solved using GMRES implementation from the PETSc library (version 3.16.3) with the relative residual tolerance set to $10^{-6}$. In addition, the problem of the pressure update (4) is solved by PCG from PETSc with the relative residual tolerance set also to $10^{-6}$.
@@ 6.1 Effect of the stopping criterion
We consider two Reynolds numbers, 100 and 300, which lead to fundamentally different regimes with respect to the dependence on time. Our main focus is the simulation of the flow at Reynolds number 300. After an initial phase, the solution is periodic, see Fig. 2 for several snapshots of vortex structures in different times and the right part of Fig. 5 for the evolution of the drag and lift coefficients in time. However, we first focus on the choice of the initial guess and the stopping criteria, and for that purpose, it is customary to perform simulations for the case of Reynolds number 100. For this lower Re, the solution is transient, and it converges to a steady state as can be deduced from the plot of the aerodynamic coefficients in Fig. 5. The vortex structure resembles the solution at the second snapshot of Fig. 2. From the linear algebraic viewpoint, the subsequent linear systems (5) are becoming closer to each other, and the solution from the preceding time step is a progressively better approximation to the solution from the actual time step. We are interested in the capability of the solver to exploit this fact and converge in a lower number of iterations.
@@
The cumulative number of PCG iterations in 4000 time steps for the different configurations is presented in Fig. 3. We consider the choice of initial guess as the vector of zeros as well as the final approximation computed in the previous time step. Two stopping criteria, (24) and (25) are compared. The cumulative number illustrates the overall demand for the number of iterations of each approach over the course of the simulation, and it provides an insight into the long-term efficiency of the criterion, here when converging to a stationary solution.
@@
Figure 3: Cumulative number of PCG iterations over all time steps for Re = 100 (transient solution, left) and Re = 300 (periodic solution, right). Iterations terminated using ${\|{\bf r}_{k}\|}/{\|{\bf r}_{0}\|}$ (denoted by ‘$\|{\bf r}_{0}\|$’) and ${\|{\bf r}_{k}\|}/{\|{\bf b}\|}$ (‘$\|{\bf b}\|$’) stopping criteria, initial guess taken as a zero vector (‘${\bf 0}$’) and as the approximate solution from the previous time step (‘${\bf x}^{\mathrm{prev}}$’).
@@
From Fig. 3, we can see that the lines corresponding to starting the solution always with the zero guess overlap (when starting with the zero guess, the stopping criteria (24) and (25) coincide as ${\bf r}_{0}={\bf b}$) and show a constant increase, indicating that the number of iterations remains almost constant in each step. When the solver for the new time step starts with the solution from the previous time step and the stopping criterion is based on the norm of the initial residual, there is a slight decrease in the number of iterations, which is approximately constant over the time steps. Even though the systems are converging to each other, the PCG method with the stopping criterion (24) based on the initial residual is not capable of reducing the number of iterations sufficiently during the simulation.
@@
Importantly, this desired behavior is obtained for the stopping criterion (25) based on the norm of the right-hand side vector, and the number of iterations decreases during the course of the time steps. As illustrated in the left part of Fig. 3, selecting an appropriate stopping criterion alone can significantly reduce the cumulative number of iterations for a simulation of the transient flow with Reynolds number 100, in our case cutting the computational time by more than 50 percent.
@@
6.2 Effect of Krylov subspace recycling
In our further experiments, we focus on the case with Re = 300, where a notable change of the regime results in a periodic behavior. In this case, as shown in the right side of Fig. 3, the choice of the initial guess and an appropriate stopping criterion alone does not lead to as substantial savings as for Re = 100, although the improvement is still considerable. In the right part of Fig. 4, we can see that in the periodic regime the residual norm of the initial guess computed from the previous approximation does not decrease. After an initial phase (approximately 1000 time steps) where the periodic solution is developed, the initial residual stagnates on a certain level (here around $0.6\times 10^{-4}$). The change in the stopping criterion then allows for a larger residual norm of the computed approximation. This softer criterion leads to a negligible difference in the resulting drag and lift coefficients; see the right panels in Fig. 5.
Figure 4: Norm of the initial and final residual over all time steps for Re = 100 (transient solution, left) and Re = 300 (periodic solution, right). Iterations terminated using ${\|{\bf r}_{k}\|}/{\|{\bf r}_{0}\|}$ (denoted as ‘$\|{\bf r}_{0}\|$’) and ${\|{\bf r}_{k}\|}/{\|{\bf b}\|}$ (‘$\|{\bf b}\|$’) stopping criteria.
Figure 5: Drag and lift coefficients over all time steps for Re = 100 (transient solution, top left) and Re = 300 (periodic solution, top right). The bottom panels give the relative difference in the coefficients when computed using the stopping criteria based on ${\|{\bf r}_{k}\|}/{\|{\bf r}_{0}\|}$ and ${\|{\bf r}_{k}\|}/{\|{\bf b}\|}$.
Let us provide some more quantitative results for Re = 300. In Table 1, we report the minimum, maximum, and mean number of PCG iterations, the mean times for the whole linear solve and for one iteration, and separately the number of iterations and time for the first step, which is excluded from the other statistics, as it includes the setup of the BDDC preconditioner. For this table (and all tables hereafter), the statistic is computed from the first 100 time steps.
@@ 6.2 Effect of Krylov subspace recycling
${\|{\bf r}_{k}\|}/{\|{\bf r}_{0}\|}$ ${\bf 0}$ 16-22(17.8) 0.50(0.028) 23 / 1.56
${\bf x}^{\mathrm{prev}}$ 16-23(17.2) 0.48(0.028) 23 / 1.61
${\|{\bf r}_{k}\|}/{\|{\bf b}\|}$ ${\bf 0}$ 16-22(17.8) 0.49(0.028) 23 / 1.37
${\bf x}^{\mathrm{prev}}$ 14-26(15.5) 0.43(0.028) 23 / 1.53
Table 1: Varying stopping criterion and initial guess, Reynolds number 300: Minimum (‘min’), maximum (‘max’) and mean (‘avg’) number of PCG iterations per time step, mean time for one step with mean time for one iteration in parentheses. For the first step (system), the number of PCG iterations and the overall time including the construction of the BDDC preconditioner are reported.
Table 1 confirms that for these simulations, the most beneficial configuration is the variant with the stopping criterion (25) based on the norm of the right-hand side vector combined with the computed approximation from the previous time step taken as the initial guess for the new time step. As expected, the mean time for one PCG iteration and the number of iterations for the first step are consistent across all four cases.
In the next experiment, we focus on the strategies for recycling the Krylov subspace. In particular, we consider the four variants of the construction of the deflation basis from Section 5.2 summarized in Table 2 together with four combinations of the stopping criteria and initial guess from the previous experiment. For these simulations, the maximal size of the deflation basis is fixed at $R=50$. In addition to the results presented in Table 1, we report the time step at which the norm of the Ritz values difference satisfied criterion (29). After this step, the deflation basis was not updated.
@@
Recycl. stop. init. #its. time [s] step 1 Ritz
strategy crit. guess min–max(avg) step(1 iter.) #its. / time [s] conv.
B1 ${\|{\bf r}_{k}\|}/{\|{\bf r}_{0}\|}$ ${\bf 0}$ 15-24(15.5) 0.47(0.030) 23 / 1.49 n/a
${\bf x}^{\mathrm{prev}}$ 13-17(15.0) 0.46(0.031) 23 / 1.51 n/a
${\|{\bf r}_{k}\|}/{\|{\bf b}\|}$ ${\bf 0}$ 14-24(15.5) 0.47(0.030) 23 / 1.71 n/a
${\bf x}^{\mathrm{prev}}$ 13-24(13.2) 0.40(0.031) 23 / 1.50 n/a
B2 ${\|{\bf r}_{k}\|}/{\|{\bf r}_{0}\|}$ ${\bf 0}$ 14-17(15.7) 0.48(0.030) 23 / 1.71 n/a
${\bf x}^{\mathrm{prev}}$ 13-17(15.4) 0.47(0.030) 23 / 1.56 n/a
${\|{\bf r}_{k}\|}/{\|{\bf b}\|}$ ${\bf 0}$ 14-17(15.7) 0.48(0.030) 23 / 1.48 n/a
${\bf x}^{\mathrm{prev}}$ 13-18(13.9) 0.43(0.031) 23 / 1.48 n/a
B3 ${\|{\bf r}_{k}\|}/{\|{\bf r}_{0}\|}$ ${\bf 0}$ 14-19(16.9) 0.52(0.030) 23 / 1.74 $>$100
${\bf x}^{\mathrm{prev}}$ 13-18(16.4) 0.50(0.030) 23 / 1.70 99
${\|{\bf r}_{k}\|}/{\|{\bf b}\|}$ ${\bf 0}$ 14-19(16.9) 0.52(0.031) 23 / 1.78 $>$100
${\bf x}^{\mathrm{prev}}$ 14-19(14.3) 0.44(0.031) 23 / 1.81 $>$100
B4 ${\|{\bf r}_{k}\|}/{\|{\bf r}_{0}\|}$ ${\bf 0}$ 12-16(13.2) 0.41(0.031) 23 / 1.81 20
${\bf x}^{\mathrm{prev}}$ 12-17(12.9) 0.39(0.031) 23 / 1.70 24
${\|{\bf r}_{k}\|}/{\|{\bf b}\|}$ ${\bf 0}$ 12-16(13.2) 0.41(0.031) 23 / 1.69 20
${\bf x}^{\mathrm{prev}}$ 11-17(11.3) 0.36(0.031) 23 / 2.19 25
@@
Table 3: Results for different recycling strategies, Reynolds number 300: Minimum (‘min’), maximum (‘max’), and mean (‘avg’) number of PCG iterations per time step, mean time for one time step with mean time for one iteration in parentheses. For the first step, the number of PCG iterations and the overall time including the construction of the BDDC preconditioner are reported. The last column (‘Ritz conv.’) gives the index of the step when the criterion (29) is met, i.e., when the Ritz values converged. The size of the deflation basis is set to $R=50$.
From Table 3 we can see that the observation made for the case with no deflation is valid also for these simulations; the most advantageous variant in terms of the number of iterations as well as the time is still the one that uses the stopping criterion (25) based on the norm of the right-hand side in combination with taking the solution from the previous time step as the initial guess for the new time step. We observe that the straightforward strategies B1 and B2 result in modest savings in both iterations and computational time with B1 being slightly more efficient.
Note that both B1 and B2 slightly outperform B3 in terms of the average number of iterations. Moreover, B1 and B2 do not require recomputing the deflation basis using (28). Due to the slow convergence of the smallest Ritz values, the deflation basis is recomputed in each time step when using B3.
@@
The best results were obtained for the strategy B4, which uses the $R$ vectors corresponding to the largest Ritz values. This choice requires, on average, the lowest number of PCG iterations, and it requires the lowest computational time for one time step. For this choice, the convergence of the Ritz values over the time steps was relatively fast, and hence the basis was recomputed only in the first few time steps. Following this experiment, we restrict ourselves to the stopping criterion based on the right-hand side vector, previous solution taken as the initial guess, and the recycling strategy B4 for further experiments.
In the next experiment, we vary the size of the deflation basis $R$. The results are presented in Table 4. As expected, the mean number of iterations decreases for increasing the size of the deflation basis. However, the time for one iteration increases with increasing $R$. In fact, the average time for the whole linear solve (i.e., all iterations) and the time step at which the Ritz values converge both reach their minima for the size of the recycling basis $R=40$, which seems to be the optimal value for this problem and setup. For $R=50$, used in the previous experiment, the performance is comparable. To provide a fair comparison, we keep the size of the deflation basis $R=50$ also for the following numerical experiments.
@@
25 12-19(12.4) 0.37(0.030) 23 / 1.86 24
30 12-18(12.3) 0.36(0.029) 23 / 1.76 26
40 11-18(11.5) 0.35(0.031) 23 / 2.02 20
50 11-17(11.3) 0.36(0.031) 23 / 2.19 25
100 9-18(10.7) 0.38(0.036) 23 / 1.80 65
200 6-18(8.7) 0.41(0.044) 23 / 1.78 89
400 5-18(8.4) 0.41(0.044) 23 / 1.94 $>$100
Table 4: Varying the maximal size of the recycling basis $R$: Minimum (‘min’), maximum (‘max’), and mean (‘avg’) number of PCG iterations per time step, mean time for one time step with mean time for one iteration in parentheses. For the first step, the number of PCG iterations and the overall time including the construction of the BDDC preconditioner are reported. The last column (‘Ritz conv.’) gives the index of the step when the criterion (29) is met.
@@ 6.3 Effect of adaptive BDDC preconditioner
In the next set of numerical experiments, we perform computations for variants of the adaptive selection of coarse degrees of freedom in the BDDC preconditioner on top of the recycling of the Krylov subspace with $R=50$. For preliminary results without recycling and with recycling using strategy B1, see our paper [11]. In the current experiment, we test several values of the prescribed target value $\tau$ from Section 3.3. For a smaller $\tau$, more eigenvectors are used in the coarse problem construction. This reduces the number of iterations, but again, each iteration gets more expensive due to a larger coarse problem. The method is combined with two types of weights in BDDC (i.e., matrix ${\bf D}_{i}$ introduced in Section 3.2), namely the scaling based on the cardinality (card) and based on the diagonal stiffness (diag).
@@
card 3.5 8-14(10.0) 0.33(0.032) 17 / 40.88 38
3.0 9-14(9.4) 0.31(0.033) 16 / 40.99 85
2.5 8-13(9.3) 0.37(0.040) 16 / 42.34 86
2.0 8-13(8.9) 0.39(0.044) 15 / 45.16 45
diag 3.5 9-12(9.7) 0.32(0.033) 17 / 41.21 53
3.0 9-12(9.2) 0.31(0.033) 17 / 42.44 49
2.5 9-12(9.2) 0.31(0.034) 16 / 42.45 34
2.0 8-12(8.5) 0.32(0.037) 16 / 42.84 43
Table 5: Varying the adaptive coarse space in BDDC by changing the threshold on eigenvalues for selecting eigenvectors for the coarse problem $\tau$: Minimum (‘min’), maximum (‘max’), and mean (‘avg’) number of PCG iterations per time step, mean time for one time step with mean time for one iteration in parentheses. The size of the deflation basis $R$ = 50. For the first step, the number of PCG iterations and the overall time including the construction of the BDDC preconditioner are reported. The last column (‘Ritz conv.’) gives the index of the step when the criterion (29) is met.
From the results in Table 5, we can conclude that the optimal threshold for eigenvalues $\tau$ is 3.0 for both types of weights, in terms of both the mean number of iterations and the mean time for the linear solver. The results indicate that the weights based on diagonal stiffness perform slightly better than using cardinality.
It is important to note that, in our current implementation, the adaptive BDDC method introduces a substantial additional cost to the preconditioner setup due to the need to solve the local eigenvalue problems (21). By comparing the average time per one time step in Tables 3 and 5, we observe that approximately 500 time steps are needed to amortize this initial cost.
Table 6 presents a summary of the results, comparing the best variant for each approach. The results show a 23% speed-up for the variant that incorporates both recycling and adaptivity compared to the one without these approaches. To further understand this behavior, Figure 6 displays the Ritz values approximating the spectrum of the preconditioned operator ${\bf M}^{-1}{\bf A}$, or ${\bf Q}{\bf M}^{-1}{\bf A}$ when deflation is used, at the 50th time step. The figure illustrates how Krylov subspace recycling and the adaptive BDDC preconditioner push the upper part of the spectrum closer to one, thereby reducing the condition number and accelerating convergence.
@@
14-26(15.5) 0.43(0.028) 23 / 41.53
$R=50$ 11-17(11.3) 0.36(0.031) 23 / 41.70
diag, $\tau=3.0$ 11-19(11.4) 0.33(0.030) 17 / 45.55
$R=50$ diag, $\tau=3.0$ 9-12(9.2) 0.31(0.033) 17 / 42.44
Table 6: Summary of the best results from Tables 15 for different acceleration strategies. Minimum (‘min’), maximum (‘max’), and mean (‘avg’) number of PCG iterations per time step, mean time for one time step with mean time for one iteration in parentheses, and the number of PCG iterations (‘step 1 #its.’), and time for the first time step. We consider the initial guess given by the computed solution to the previous system and the stopping criterion (25).
@@
To investigate the scalability and efficiency of the acceleration strategies studied for the smaller mesh size, we finally consider a larger computational mesh, consisting of 15.6 million unknowns for pressure solved on 4096 subdomains and processor cores. We monitor the same metrics of iterations and times as for the smaller problem. Benchmark simulations are carried out again in four different modes: without Krylov subspace recycling, with recycling, with adaptive BDDC, and with a combination of both Krylov subspace recycling and adaptive BDDC. The results are summarized in Table 7.
variant stop. init. #its. time [s] step 1 Ritz
recycl. adapt. crit. guess min–max(avg) step(1 iter.) #its. / time [s] conv.
$\frac{\|{\bf r}_{k}\|}{\|{\bf r}_{0}\|}$ ${\bf 0}$ 34-47(36.7) 4.03(0.109) 47 / 8.55 n/a
${\bf x}^{\mathrm{prev}}$ 34-47(36.0) 3.98(0.110) 47 / 8.37 n/a
$\frac{\|{\bf r}_{k}\|}{\|{\bf b}\|}$ ${\bf 0}$ 34-47(36.7) 3.88(0.106) 47 / 8.85 n/a
${\bf x}^{\mathrm{prev}}$ 31-54(33.3) 3.59(0.107) 47 / 8.24 n/a
$R=50$ $\frac{\|{\bf r}_{k}\|}{\|{\bf b}\|}$ ${\bf x}^{\mathrm{prev}}$ 22-35(23.3) 2.67(0.110) 47 / 8.43 18
$\tau=3.0$ 17-30(18.4) 2.31(0.125) 26 / 179.86 n/a
$R=50$ $\tau=3.0$ 13-21(14.3) 1.89(0.132) 26 / 177.87 17
Table 7: Larger computational mesh with 15.6 million pressure unknowns: analogy to the results from Tables 1 and 6.
The trends observed for the smaller problem are confirmed for the larger problem; see Table 7. Using the stopping criterion based on the right-hand side vector and initializing the linear solver at each time step with the solution from the previous step results in both the lowest number of iterations and the shortest average time for solving the linear system. When recycling, adaptivity, and their combination are introduced, the results align with those from the smaller problem, but with the differences being more pronounced. The combination of adaptive BDDC and Krylov subspace recycling proves to be the most effective in reducing the iteration count and, more importantly, the computational time for the linear solver. In particular, these acceleration techniques lead to reducing the computational time approximately by 50 %.
Let us also note for completeness that we have performed experiments with different time step sizes, namely, we tested 10$\times$ larger and 10$\times$ smaller time steps as well. This change had negligible effect on the convergence of the studied Poisson problem of the pressure corrector, and hence the conclusions of the previous experiments remain valid.

### v3

@@ 6 Results
In this section, we present numerical results for the simulation of incompressible flow around a unit sphere. We focus on multiple variants of the three-level BDDC method for the Poisson problem for pressure corrector (3) with a fixed matrix and time-dependent right-hand side vector. The computational mesh of Taylor–Hood hexahedral elements leads to 1.4M unknowns for pressure, and it is decomposed into 1024 subdomains; see Fig. 1. The time step is constant and set to 0.05 s.
@@ 6.1 Effect of the stopping criterion
Our main focus is the simulation of the flow at Reynolds number 300. After an initial phase, the solution is periodic, see Fig. 2 for several snapshots of vortex structures in different times and the right part of Fig. 5 for the evolution of the drag and lift coefficients in time. In this experiment, we first focus on the effect of the choice of the initial guess and the stopping criteria, and for that purpose, it is customary to perform also simulations for the case of Reynolds number 100. For this lower Re, the solution is transient and converges to a steady state as can be deduced from the plot of the aerodynamic coefficients in Fig. 5 (left). The vortex structure resembles the solution at the second snapshot of Fig. 2. From the linear algebraic viewpoint, the right-hand sides in subsequent linear systems (5) are becoming closer to each other, and the solution from the preceding time step is a progressively better approximation to the solution from the actual time step; recall that the system matrix and preconditioner remain the same in all time steps. We are interested in the capability of the solver to exploit this fact and converge in a lower number of iterations. The multilevel BDDC preconditioner uses the inverse cardinality (card) as the interface weights.
@@
The cumulative number of PCG iterations in 4000 time steps for the different configurations is presented in Fig. 3, left. We consider the choice of initial guess as the vector of zeros as well as the final approximation computed in the previous time step. Two stopping criteria, (23) and (24) are compared, both with $tol=10^{-6}$. The cumulative number illustrates the overall demand for the number of iterations of each approach over the course of the simulation.
@@
Figure 3: Cumulative number of PCG iterations over all time steps for Re = 100 (transient solution, left) and Re = 300 (periodic solution, right). Iterations terminated using ${\|{\bf r}_{k}\|}/{\|{\bf r}_{0}\|}$ (denoted by $\|{\bf r}_{0}\|$’) and ${\|{\bf r}_{k}\|}/{\|{\bf b}\|}$ ($\|{\bf b}\|$’) stopping criteria, initial guess taken as a zero vector ($\bf 0$}’) and as the approximate solution from the previous time step (${\bf x}^{\mathrm{prev}}$’).
@@
From Fig. 3, we can see that the lines corresponding to starting the solution always with the zero guess overlap (when starting with the zero guess, the stopping criteria (23) and (24) coincide as ${\bf r}_{0}={\bf b}$) and show a constant increase, indicating that the number of iterations remains almost constant in each step. When the solver starts with the solution from the previous time step and the stopping criterion is based on the norm of the initial residual, there is a marginal decrease in the number of iterations. The desired behavior is obtained only for the stopping criterion (24) based on the norm of the right-hand side vector, and the number of iterations decreases during the course of the time steps.
@@
As illustrated in the left part of Fig. 3, selecting the stopping criterion (24) can significantly reduce the cumulative number of iterations for a simulation of the transient flow with Reynolds number 100, in our case cutting the computational time by more than 50 percent.
@@
In our further experiments, we focus on the case with Re = 300, where a notable change of the regime results in a periodic behavior. In this case, as shown in the right side of Fig. 3, the choice of the initial guess and an appropriate stopping criterion alone does not lead to as substantial savings as for Re = 100, although the improvement is still considerable. In the right part of Fig. 4, we can see that in the periodic regime the residual norm of the initial guess computed from the previous approximation does not decrease. After an initial phase (approximately 1000 time steps) where the periodic solution is developed, the initial residual stagnates on a certain level (here around $6\times 10^{-5}$). The change in the stopping criterion then allows for a larger residual norm of the computed approximation. This softer criterion leads to a negligible difference in the resulting drag and lift coefficients; see the right panel in Fig. 5.
Figure 4: Norm of the initial and final residual over all time steps for Re = 100 (transient solution, left) and Re = 300 (periodic solution, right). Iterations terminated using ${\|{\bf r}_{k}\|}/{\|{\bf r}_{0}\|}$ (denoted as $\|{\bf r}_{0}\|$’) and ${\|{\bf r}_{k}\|}/{\|{\bf b}\|}$ ($\|{\bf b}\|$’) stopping criteria. The time from which the statistics are computed is marked by a vertical line for Re = 300.
Figure 5: Drag and lift coefficients over all time steps for Re = 100 (transient solution, left) and Re = 300 (periodic solution, right). The lines for different stopping criteria visually overlap, with relative difference in most of the iterations below $10^{-6}$.
Let us provide some more quantitative results for Re = 300. In Table 1, we report the minimum, maximum, and mean number of PCG iterations, the mean times for the whole linear solve and for one iteration, and separately the number of iterations and time for the first step, which is excluded from the other statistics, as it includes the setup of the BDDC preconditioner. For this table (and all tables hereafter), the statistics is computed from the time steps 2001 to 2100, i.e., after the periodic solution is developed, see Fig. 4.
@@
${\|{\bf r}_{k}\|}/{\|{\bf r}_{0}\|}$ $\bf 0$} 20-20(20) 0.27(0.014) 20 / 0.79
${\bf x}^{\mathrm{prev}}$ 19-20(20.0) 0.28(0.014) 20 / 0.79
${\|{\bf r}_{k}\|}/{\|{\bf b}\|}$ $\bf 0$} 20-20(20) 0.27(0.014) 20 / 0.76
${\bf x}^{\mathrm{prev}}$ 17-17(17) 0.24(0.014) 20 / 0.77
Table 1: Varying stopping criterion and initial guess, Re = 300: Minimum (‘min’), maximum (‘max’), and mean (‘avg’) number of PCG iterations per time step, mean time for one step with mean time for one iteration in parentheses. For the first step (system), the number of PCG iterations and the overall time including the construction of the BDDC preconditioner are reported.
Table 1 confirms that for these simulations, the most beneficial configuration is the variant with the stopping criterion (24) based on the norm of the right-hand side vector combined with the computed approximation from the previous time step taken as the initial guess for the new time step. We therefore consider this choice of the initial guess and the stopping criterion in all experiments that follow. As expected, the mean time for one PCG iteration and the number of iterations for the first step are consistent across all four cases.
6.2 Effect of Krylov subspace recycling
In the next experiment, we focus on the strategies for recycling the Krylov subspace. In particular, we consider four variants of the construction of the deflation basis from Section 4.2 summarized in Table 2. For these simulations, the maximal size of the deflation basis is fixed at $R=50$. In addition to the results presented in Table 1, we report the time step at which the norm of the Ritz values difference satisfied criterion (22). After this step, the deflation basis was not updated. Recall that the statistics is computed from the time steps 2001 to 2100.
@@ 6.2 Effect of Krylov subspace recycling
Recycl. #its. time [s] step 1 Ritz
strategy min–max(avg) step(1 iter.) #its. / time [s] conv.
B1 14-26(15.1) 0.24(0.016) 20 / 0.82 n/a
B2 12-16(13.6) 0.23(0.016) 20 / 0.84 n/a
B3 13-16(14.4) 0.24(0.016) 20 / 1.76 >>100
B4 13-16(13.4) 0.22(0.016) 20 / 1.70 46
@@
Table 3: Results for different recycling strategies, Re = 300: Minimum (‘min’), maximum (‘max’), and mean (‘avg’) number of PCG iterations per time step, mean time for one time step with mean time for one PCG iteration in parentheses. For the first step, the number of PCG iterations and the overall time including the construction of the BDDC preconditioner are reported. The last column (‘Ritz conv.’) gives the index of the step when the criterion (22) is met, i.e., when the Ritz values converged. The size of the deflation basis is set to $R=50$.
From Table 3 we observe that all strategies reduce the number of PCG iterations, which was originally 17. The time of single PCG iteration is the same, up to two valid digits, for all the recycling strategies, and it increased by 14 % in comparison with the variant without deflation. Note that B1 and B2 do not require recomputing the deflation basis using (21), but the cost of recomputing the basis is negligible with respect to other operations. Overall, the savings in computational time per step are not very satisfactory for B1 and B3. Due to the slow convergence of the smallest Ritz values, the deflation basis is recomputed in each time step when using B3.
@@
The best results were obtained for the strategy B4, which uses the $R$ vectors corresponding to the largest Ritz values. This choice requires, on average, the lowest number of PCG iterations, and the lowest computational time for one time step. In this setting, also the strategy B2, which is a bit simpler to implement, can also be recommended. Following this experiment, we restrict ourselves to the recycling strategy B4 for further experiments.
In the next experiment, we vary the size of the deflation basis $R$. The results are presented in Table 4. As expected, increasing the size of the deflation basis $R$ reduces the mean number of iterations, but increases the time required for each of them. Choosing a proper size of the deflation basis can be a difficult task and, in practice, this might be done based on memory limitations, i.e., choosing the size as the maximal number of the vectors that can be stored. This experiment reveals, however, that similar timings can be obtained for a range of sizes. The choice $R=50$, used in the previous experiment, is one with nearly the best computational time for a single time step. To provide a fair comparison, we keep the size of the deflation basis $R=50$ also for the following numerical experiments.
@@
25 16-17(16.2) 0.24(0.015) 20 / 1.88 12
30 13-16(13.9) 0.22(0.016) 20 / 1.75 15
35 13-16(13.2) 0.21(0.016) 20 / 1.81 17
50 13-16(13.4) 0.22(0.016) 20 / 1.70 46
100 9-16(11.7) 0.24(0.020) 20 / 1.78 >>100
200 8-16(10.2) 0.23(0.022) 20 / 1.97 >>100
400 8-16(10.3) 0.23(0.022) 20 / 2.07 >>100
Table 4: Varying the maximal size of the recycling basis $R$ in recycling strategy B4: Minimum (‘min’), maximum (‘max’), and mean (‘avg’) number of PCG iterations per time step, mean time for one time step with mean time for one PCG iteration in parentheses. For the first step, the number of PCG iterations and the overall time including the construction of the BDDC preconditioner are reported. The last column (‘Ritz conv.’) gives the index of the step when the criterion (22) is met.
@@ 6.3 Effect of adaptive BDDC preconditioner
In the next set of numerical experiments, we perform computations for variants of the adaptive selection of coarse degrees of freedom in the BDDC preconditioner on top of the recycling of the Krylov subspace with $R=50$. For preliminary results without recycling and with recycling using strategy B1, see our paper [11]. In the current experiment, we test several values of the prescribed target value $\tau$ described in Section 3.3. For a smaller $\tau$, more eigenvectors are used in the construction of the coarse problem. This reduces the number of iterations, but again, each iteration gets more expensive due to a larger coarse problem. The method is combined with two types of weights in BDDC (i.e., matrices ${\bf D}_{i}$ introduced in Section 3.2), namely the scaling based on the cardinality (card) and based on the diagonal stiffness (diag).
@@
card 3.5 9-12(11.3) 0.20(0.017) 14 / 17.85 73
3.0 8-12(10.3) 0.19(0.018) 14 / 18.23 74
2.5 8-11(10.0) 0.19(0.019) 14 / 18.85 69
2.0 8-11(8.7) 0.25(0.025) 14 / 21.69 76
diag 3.5 10-12(11.0) 0.19(0.017) 15 / 18.14 33
3.0 10-12(10.2) 0.18(0.018) 15 / 18.41 39
2.5 10-12(10.1) 0.19(0.019) 15 / 19.78 33
2.0 9-12(10.0) 0.20(0.020) 16 / 19.80 30
Table 5: Varying the adaptive coarse space in BDDC by changing the threshold on eigenvalues for selecting eigenvectors for the coarse problem $\tau$: Minimum (‘min’), maximum (‘max’), and mean (‘avg’) number of PCG iterations per time step, mean time for one time step with mean time for one PCG iteration in parentheses. The size of the deflation basis $R$ = 50. For the first step, the number of PCG iterations and the overall time including the construction of the BDDC preconditioner are reported. The last column (‘Ritz conv.’) gives the index of the step when the criterion (22) is met.
From the results in Table 5, we can conclude that the optimal threshold for eigenvalues $\tau$ is 3.0 for both types of weights, in terms of the mean time for the linear solver. The results indicate that the weights based on diagonal stiffness (diag) perform slightly better than those based on cardinality (card). Consequently, diag weights are used in further experiments of this paper.
It is important to note that in our current implementation, the adaptive BDDC method introduces a substantial additional cost to the preconditioner setup, as it requires solving local eigenvalue problems (16). By comparing the average time per one time step in Tables 3 and 5, we observe that approximately 350 time steps are needed to amortize this initial cost.
Table 6 presents a summary of the results, comparing the best variant for each approach. The results show a 25% speed-up in computational time for one time step for the variant that incorporates both recycling and adaptivity compared to the one without these approaches. To further understand this behavior, Figure 6 displays the Ritz values approximating the spectrum of the preconditioned operator ${\bf M}^{-1}{\bf A}$, or ${\bf Q}{\bf M}^{-1}{\bf A}$ when deflation is used. The figure illustrates how Krylov subspace recycling and the adaptive BDDC preconditioner push the upper part of the spectrum closer to one, thereby reducing the condition number and accelerating convergence.
@@
17-17(17) 0.24(0.014) 20 / 40.77
$R=50$ 13-16(13.4) 0.22(0.016) 20 / 41.70
diag, $\tau=3.0$ 12-13(12.8) 0.19(0.015) 15 / 18.79
$R=50$ diag, $\tau=3.0$ 10-12(10.2) 0.18(0.018) 15 / 18.41
Table 6: Summary of the best results from Tables 15 for different acceleration strategies. Minimum (‘min’), maximum (‘max’), and mean (‘avg’) number of PCG iterations per time step, mean time for one time step with mean time for one PCG iteration in parentheses, and the number of PCG iterations (‘step 1 #its.’), and time for the first time step. We consider the initial guess given by the computed solution to the previous system and the stopping criterion (24).
@@
To investigate the efficiency of the acceleration strategies studied for the smaller mesh size, we finally consider a larger computational mesh, consisting of 15.6 million unknowns for pressure solved on 4096 subdomains and processor cores. We monitor the same metrics of iterations and times as for the smaller problem. Benchmark simulations are carried out again in four different modes: without Krylov subspace recycling, with recycling, with adaptive BDDC, and with a combination of both Krylov subspace recycling and adaptive BDDC. The results are summarized in Table 7.
variant #its. time [s] step 1
recycling adaptivity min–max(avg) step(1 iter.) #its. / time [s]
37-38(37.7) 3.94(0.104) 44 / 47.00
$R=50$ 26-30(26.1) 2.93(0.113) 44 / 48.17
diag, $\tau=3.0$ 20-22(20.6) 2.53(0.122) 24 / 114.06
$R=50$ diag, $\tau=3.0$ 16-19(16.8) 2.21(0.132) 24 / 113.62
Table 7: Larger computational mesh with 15.6 million pressure unknowns and 4096 subdomains: analogy to the results from Table 6.
The trends observed for the smaller problem are confirmed for the larger problem (compare Table 6 and Table 7). When recycling, adaptivity, and their combination are introduced, the results align with those from the smaller problem, but with the differences being even more pronounced. The combination of adaptive BDDC and Krylov subspace recycling proves to be the most effective in reducing the iteration count and, more importantly, the computational time for the linear solver. In particular, these acceleration techniques lead to reducing the computational time per one time step by more than 40 %.
When comparing the times reported in Tables 6 and 7, we observe a noticeable increase for the larger problem. This behavior is partly in line with the design of our experiments: since the second problem employs larger subdomains, weak scalability is not expected. Furthermore, by constraining the size of the local problems for the velocity unknowns, the corresponding local problems for the pressure corrector become very small, and hence communication overhead dominates the computational time.
[…]